OpenStudy (anonymous):

evaluate the expression A=-1-|-1-[-1-(-1)-1]-1|-1

OpenStudy (anonymous):

[-1-(-1)-1] -1+1-1 -1 []=-1 -1-|-1-[-1]-1|-1 |-1-[-1]-1| |-1| -1-|-1|-1 -1-1-1=-3